# Panning for Gold: Expanding Domain-Specific Knowledge Graphs with General Knowledge

### v1

@@ Abstract.
Domain-specific knowledge graphs (DKGs) are essential for supporting intelligent applications in specialized fields, yet they often suffer from limited coverage and incompleteness compared to general knowledge graphs (GKGs) such as Wikipedia and YAGO. Existing tasks to enrich DKGs rely primarily on extracting knowledge from unstructured data or completing KGs through internal reasoning, but the scope and quality of such integration remain limited. This highlights a critical gap: little systematic exploration has been conducted on how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. To address this gap, we propose a new and practical task: domain-specific knowledge graph fusion (DKGF), which aims to mine and integrate relevant facts from general knowledge graphs into domain-specific knowledge graphs to enhance their completeness and utility. Unlike previous research, this new task faces two key challenges: ① high ambiguity of domain relevance, i.e., difficulty in determining whether knowledge from a GKG is truly relevant to the target domain, and ② cross-domain knowledge granularity misalignment, i.e., GKG facts are typically abstract and coarse-grained, whereas DKGs frequently require more contextualized, fine-grained representations aligned with particular domain scenarios.
To tackle these challenges, we propose ExeFuse, a simple yet effective Fact-as-Program paradigm that reformulates DKGF as executable semantic reasoning over DKGs. Specifically, ExeFuse interprets each GKG fact as a latent semantic program, maps abstract relations to granularity-aware operators, and verifies domain relevance through program executability on the target DKG. By unifying relevance assessment and granularity transformation within a single probabilistic framework, ExeFuse enables precise identification and integration of domain-relevant, consistent knowledge from GKGs. We construct two new benchmark datasets (DKGF(W-I) and DKGF(Y-I)) and propose 21 representative benchmark configurations to systematically assess DKGF performance. Extensive experiments highlight the value of the new task and demonstrate the effectiveness of ExeFuse, providing the first standardized evaluation suite for this emerging task. The source codes and datasets are available at https://github.com/eduzrh/DKGF_benchmark.
@@ 1. Introduction
Domain-specific knowledge graphs (DKGs) have emerged as essential infrastructures for representing structured knowledge in specialized domains and enabling intelligent applications (Cao et al., 2024; Ma et al., 2024; Huang et al., 2023, 2024; Wang et al., 2024; Tian et al., 2024; Geisler et al., 2025; Zhu et al., 2023). For example, ICEWS [^1] https://www.andybeger.com/icews captures political events for international relations analysis, while PubChem [^2] https://pubchem.ncbi.nlm.nih.gov provides chemical knowledge widely used in drug discovery. These DKGs demonstrate the great potential of domain-specific structured knowledge in powering real-world applications (Chen et al., 2024d; Chai et al., 2023; Huo et al., 2024; Zhang et al., 2024c, a; Khan, 2023). Nevertheless, compared with well-maintained general knowledge graphs (GKGs) such as Wikipedia (Erxleben et al., 2014) and YAGO (Suchanek et al., 2007), which benefit from community-driven contributions and timely updates, DKGs often suffer from limited completeness. They inherently face challenges due to restricted data sources (Lin et al., 2023), domain-specific construction methods (Khan et al., 2025; Kondylakis et al., 2025; Khan, 2023), and costly maintenance efforts (Akrami et al., 2020; Abu-Salih, 2021; Li et al., 2020; Chen et al., 2024a). These limitations create a substantial barrier for DKGs to fully support downstream applications in specialized domains (Khan et al., 2024; Le et al., 2024). Therefore, a pressing challenge is to efficiently and accurately enrich DKGs, improving their coverage and the ability to support domain-specific applications.
@@
Although currently there is no work that directly investigates how to enrich DKGs, adapting existing techniques developed for supplementing GKGs may be a viable solution. Specifically, these efforts can be broadly divided into two categories: integration from unstructured knowledge and from structured knowledge. ① The former category mainly includes manual annotation, entity linking, and relation triple extraction (Jin et al., 2021; Nie et al., 2024; Li et al., 2024b). However, as discussed earlier, domain-specific unstructured data sources are inherently limited in scope and coverage, constraining the effectiveness of these text-based extraction methods. Moreover, the specialized terminology and complex linguistic patterns in domain-specific texts further exacerbate the challenges of accurate knowledge extraction. ② The latter category focuses on knowledge graph completion (Wei et al., 2024), which predicts missing triples through internal inference within the existing KG. While these solutions can discover implicit relationships, the knowledge they generate is fundamentally bounded by the original KG’s scope and cannot introduce genuinely new external knowledge. Given the limitations of this existing research and the rich, continuously updated knowledge available in comprehensive GKGs like Wikipedia and YAGO, a natural question arises: Can we systematically leverage the abundant knowledge in GKGs to supplement and enrich DKGs?
@@
To this end, we propose a new and important task setting, i.e., domain-specific knowledge graph fusion (DKGF), which aims to mine the highly relevant knowledge in the general knowledge graph to enhance the knowledge coverage of domain-specific knowledge graph, as illustrated in Figure 1. Unlike the widely studied knowledge graph alignment task (Leone et al., 2022; Zeng et al., 2022; Tian et al., 2024), which focuses on identifying entities that are simultaneously present and semantically equivalent across different KGs, the DKGF task essentially focuses on how to extract knowledge from a large-scale semantically complex GKG that is closely related to the target domain but have not yet been included in the DKG, and then effectively integrate them, while maintaining the semantic consistency and structural integrity of the DKG itself. As a result, the DKGF task poses new challenges:
@@
Challenge I: High Ambiguity of Domain Relevance. Firstly, GKGs contain a large amount of knowledge with low similarity to that in DKGs, yet such information may still provide critical factual associations for enriching DKGs in reality. As illustrated in Figure 1, during the mining process of relevant knowledge from GKGs, the academic institution “Brookings Institution” from Washington D.C. lacks explicit connections to entities in the political crisis event DKG (such as “Iran”), resulting in low semantic similarity between entities. Nonetheless, in reality, “Brookings Institution” maintains many important connections with DKG entities and is highly likely to constitute factual associations with them. This discrepancy creates ambiguity in determining whether knowledge from GKGs is truly domain-relevant, making it difficult to directly apply traditional graph matching methods or language models (Pan et al., 2024a) to this new task.
@@
To address the above challenges, we propose a new and effective framework, ExeFuse, which introduces a Fact-as-Program paradigm that reformulates DKGF as executable semantic reasoning over DKG. The core insight is that domain relevance can be operationalized as program executability: a general fact is domain-relevant if and only if it can be “executed” on the DKG structure to generate valid domain facts. Specifically, ExeFuse interprets each GKG fact as a latent semantic program, maps abstract relations to granularity-aware operators through a three-level hierarchy (coarse, medium, fine), and verifies domain relevance by assessing whether the composed program can successfully execute on the target DKG. By unifying relevance assessment and granularity transformation within a single probabilistic framework, ExeFuse enables precise identification and integration of domain-relevant, fine-grained knowledge from GKGs while maintaining semantic consistency and structural integrity of the target DKG.
Notably, to fill in the gap of DKGF benchmarks, we construct two new datasets: DKGF(W-I) and DKGF(Y-I), and offer 21 benchmark configurations for systematic evaluation. Extensive experiments highlight the value of the new task and demonstrate the effectiveness of ExeFuse, providing the first standardized evaluation suite for this emerging task.
@@
    A New and Important Research Task. To the best of our knowledge, this is the first work that investigates how to enrich DKG with facts from GKG. We formally formulate this problem as the domain-specific knowledge graph fusion (DKGF) problem, and identify the key challenges brought by this new task, i.e., high ambiguity in domain relevance and cross-domain knowledge granularity misalignment, establishing a foundational framework and benchmarks to facilitate future research in this practical yet underexplored area.
@@
    A Simple and Effective DKGF Paradigm. We propose ExeFuse, a new Fact-as-Program paradigm that reformulates fusion as executable semantic reasoning, effectively addressing both high ambiguity of domain relevance and cross-domain granularity misalignment through a unified probabilistic framework.
@@
    Comprehensive Validation Through Extensive Experiments. To address the scarcity of standardized benchmarks for DKGF, we develop two new benchmark datasets, DKGF(W-I) and DKGF(Y-I), specifically tailored to this new task. Furthermore, we propose 21 representative and classic benchmark configurations for systematic evaluation. Extensive experiments highlight the value of the new task and demonstrate the effectiveness of ExeFuse, providing the first standardized evaluation suite for this emerging task.
@@ 2. Related work
To the best of our knowledge, DKGF represents a new task in domain-specific data integration. There are some closely related research efforts that can also be applied to enriching DKGs, which can be broadly grouped into two categories: integration from structured knowledge and integration from unstructured knowledge.
2.1. Integration from Structured Knowledge
2.1.1. Entity alignment
Entity alignment aims to identify equivalent entities in different knowledge graphs. Existing approaches include Translation-based (Zhao et al., 2025b; Sun et al., 2018; Bordes et al., 2013; Fanourakis et al., 2024), GNN-based (Wang et al., 2018; Chen et al., 2022b; Mao et al., 2021; Xu et al., 2021; Cai et al., 2022; Liu et al., 2023) and other methods (Wang et al., 2023a; Jiang et al., 2024c; Tang et al., 2020; Jiang et al., 2024b; Chen et al., 2024b; Yang et al., 2024; Zhao et al., 2025a). The early EA methods are primarily translation-based methods, such as MTransE (Chen et al., 2017), BootEA (Sun et al., 2018), and AlignE (Sun et al., 2018). These methods leverage the TransE framework (Bordes et al., 2013) for knowledge representation, but face limitations in handling complex structural and semantic contexts. With the development of deep graph neural networks, GNN-based methods have received widespread attention, including GCN-Align (Wang et al., 2018), RDGCN (Chen et al., 2022b), and Dual-AMN (Mao et al., 2021), which utilize graph convolutional networks to capture neighborhood structures, with temporal extensions like TEA-GNN (Xu et al., 2021), STEA (Cai et al., 2022), Dual-Match (Liu et al., 2023), and MGTEA (Zeng et al., 2024). Other approaches, such as FuAlign (Wang et al., 2023a), Simple-HHEA (Jiang et al., 2024c), and BERT-INT (Tang et al., 2020), address knowledge graph heterogeneity by integrating side information and pre-trained language models. Recently, LLM-based methods like LLM4EA (Chen et al., 2024b), ChatEA (Jiang et al., 2024b), and AdaCoAgentEA (Zhao et al., 2025a) have gained attention due to their few-shot learning capabilities and enhanced semantic understanding.
2.1.2. Knowledge Graph Completion
Knowledge graph completion aims to discover latent triples within existing knowledge structures (Wang et al., 2017). This task encompasses multiple subtasks, including triple classification (Bordes et al., 2013) and entity prediction (Bordes et al., 2013), all of which share the common goal of developing robust mechanisms to assess triple plausibility. Contemporary methods can be broadly divided into two categories: embedding-based techniques and pre-trained language model-based (PLM-based) methods. Embedding-based approaches (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016; Sun et al., 2019) focus on projecting entities and relations from knowledge graphs into continuous vector spaces. These methods leverage structural patterns within knowledge graphs to model triple credibility through carefully crafted scoring functions, enabling self-supervised learning of entity and relation representations. In contrast, PLM-based methods reframe knowledge graph completion as text processing tasks by adapting pre-trained language models (Devlin et al., 2019b) for knowledge graph scenarios. These approaches transform brief textual descriptions into structured input sequences for PLM encoding. Pioneering this direction, KG-BERT (Yao et al., 2019a) introduced the paradigm of treating KGC as binary text classification. Subsequent developments, including MTL-KGC (Kim et al., 2020) and StAR (Wang et al., 2021), enhanced this foundation by incorporating additional training objectives such as relation classification and triple ranking, alongside more sophisticated triple encoding strategies. PKGC (Lv et al., 2022a) advanced the field by employing manually designed prompt templates to capture semantic relationships within triples. Furthermore, generative approaches exemplified by KGT5 (Saxena et al., 2022; Chen et al., 2022a) have pioneered sequence-to-sequence paradigms for generative KGC (Ye et al., 2022b), utilizing encoder-decoder architectures such as T5 (Raffel et al., 2020).
2.2. Integration from Unstructured Knowledge
2.2.1. Relation Triple Extraction
Relation triple extraction aims to extract structured triplets from unstructured text (Hu et al., 2023; Li et al., 2024a). Early approaches focus on supervised (Wang and Lu, 2020; Ye et al., 2022a) and semi-supervised methods (Lin et al., 2019) that require sufficient labeled data. However, data scarcity in real-world scenarios limits model performance on uncommon relations, leading to increased focus on zero-shot relation triple extraction (Chia et al., 2022; Li et al., 2025b), which extracts unseen relations and corresponding entities without prior training examples. Current zero-shot relation triple extraction methods can be categorized into two main approaches. Data augmentation-based methods leverage synthetic data to improve generalization, exemplified by RelationPrompt (Chia et al., 2022), which employs pre-trained BART (Lewis et al., 2020) and GPT-2-generated synthetic data (Radford et al., 2019), with subsequent works incorporating external knowledge to enhance synthetic data quality (Guo et al., 2024; Gong and Eldardiry, 2024). Knowledge-driven methods emphasize leveraging prior knowledge from pre-trained models, such as ZETT (Kim et al., 2023), which reformulates zero-shot relation triple extraction as a template completion task to intuitively adapt to new relations. Recent advances like TGM (Li et al., 2025a) address the limitation of insufficient generalization optimization during training by adopting a bilevel optimization perspective to enhance model generalization performance.
2.2.2. Entity Linking
Entity linking aims to identify entity mentions in text and link them to corresponding entities in knowledge graphs. Traditional approaches (Moussallem et al., 2017; Jiang et al., 2024a; Ding et al., 2024; Graciotti et al., 2025; Vollmers et al., 2025) follow a two-phase pipeline involving span detection and entity disambiguation. During span detection phase, prevalent methods utilize named entity recognition techniques to locate pertinent text spans. In the disambiguation phase, potential entities are retrieved from knowledge graphs and matched to the most suitable candidates through pre-constructed search indexes, with each index entry representing a KG entity. Early disambiguation methods like MAG (Moussallem et al., 2017) and DoSeR (Zwicklbauer et al., 2016) rely on pre-built indices with surface forms and contextual information, while embedding-based approaches such as BLINK (Wu et al., 2020) employ bi-encoder models for candidate generation and KNN-search for entity matching. Recent end-to-end methods bypass the separate NER step and directly perform entity annotation, including sequence-to-sequence approaches (Kolitsas et al., 2018; van Hulst et al., 2020) and fine-tuned language models like GENRE (Cao et al., 2021) that leverage autoregressive generation and constrained decoding for improved linking performance. Furthermore, some methods (Vollmers et al., 2025; Graciotti et al., 2025; Ding et al., 2024) introduce LLM-based augmentation to improve the entire entity linking pipeline.
@@ 2.2.2. Entity Linking
Despite these efforts, existing research predominantly relies on extracting knowledge from unstructured data or completing KGs through internal reasoning, while the scope and quality of knowledge integration remain significantly limited. This situation exposes a critical gap: There has been little systematic exploration of how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. DKGF is therefore an essential and indispensable task. Furthermore, this new task differs fundamentally from existing ones. Entity alignment outputs same entity pairs across KGs, whereas DKGF identifies missing but relevant entities and produces new facts. Knowledge graph completion focuses on reasoning within a single KG, while DKGF integrates knowledge from external GKG. Relation triple extraction operates on unstructured text, whereas DKGF leverages structured GKG. Entity linking resolves text-to-KG disambiguation, whereas DKGF enhances DKG coverage via cross structured graph discovery and fusion. Thus, existing methods cannot be directly applied to this new task, which faces the greater challenges of high ambiguity of domain relevance and cross-domain knowledge granularity misalignment.

### v2

@@ Abstract.
Domain-specific knowledge graphs (DKGs) are critical yet often suffer from limited coverage compared to General Knowledge Graphs (GKGs). Existing tasks to enrich DKGs rely primarily on extracting knowledge from external unstructured data or completing KGs through internal reasoning, but the scope and quality of such integration remain limited. This highlights a critical gap: little systematic exploration has been conducted on how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. To address this gap, we propose a new and practical task: domain-specific knowledge graph fusion (DKGF), which aims to mine and integrate relevant facts from general knowledge graphs into domain-specific knowledge graphs to enhance their completeness and utility. Unlike previous research, this new task faces two key challenges: ① high ambiguity of domain relevance, i.e., difficulty in determining whether knowledge from a GKG is truly relevant to the target domain, and ② cross-domain knowledge granularity misalignment, i.e., GKG facts are typically abstract and coarse-grained, whereas DKGs frequently require more contextualized, fine-grained representations aligned with particular domain scenarios. To address these, we present ExeFuse, a neuro-symbolic framework based on a novel Fact-as-Program paradigm. ExeFuse treats fusion as an executable process, utilizing neuro-symbolic execution to infer logical relevance beyond surface similarity and employing target space grounding to calibrate granularity. We construct two new datasets to establish the first standardized evaluation suite for this task. Extensive experiments demonstrate that ExeFuse effectively overcomes domain barriers to achieve superior fusion performance. The source codes and datasets are available at https://github.com/eduzrh/code_datasets.
@@ 1. Introduction
Domain-specific knowledge graphs (DKGs) have emerged as essential infrastructures for representing structured knowledge in specialized domains and enabling intelligent applications (Zhang et al., 2025; Kejriwal, 2019; Cao et al., 2024; Geisler et al., 2025; Hassani-Pak et al., 2020; Faghihi et al., 2021; Gupta et al., 2024). For example, ICEWS [^1] https://www.andybeger.com/icews captures political events for international relations analysis, while PubChem [^2] https://pubchem.ncbi.nlm.nih.gov provides chemical knowledge widely used in drug discovery. These DKGs demonstrate the great potential of domain-specific structured knowledge in powering real-world applications (Ji et al., 2022; Chen et al., 2024d; Huo et al., 2024; Zhang et al., 2024c, a). Nevertheless, compared with well-maintained general knowledge graphs (GKGs) such as Wikipedia (Erxleben et al., 2014) and YAGO (Suchanek et al., 2007), which benefit from community-driven contributions and timely updates, DKGs often suffer from limited completeness. They inherently face challenges due to restricted data sources (Lin et al., 2023), domain-specific construction methods (Khan et al., 2025; Kondylakis et al., 2025), and costly maintenance efforts (Yang et al., 2025a; Abu-Salih, 2021; Li et al., 2020; Chen et al., 2024a). These limitations create a substantial barrier for DKGs to fully support downstream applications in specialized domains (Khan et al., 2024; Paulheim, 2017). Therefore, a pressing challenge is to efficiently and accurately enrich DKGs, improving their coverage and the ability to support domain-specific applications.
@@
Although currently there is no work that directly investigates how to enrich DKGs, adapting existing techniques developed for supplementing GKGs may be a viable solution. Specifically, these efforts can be broadly divided into two categories: extraction from external unstructured sources and reasoning based on internal information. ① The former category aims to acquire new knowledge from external unstructured texts, mainly including manual annotation, entity linking, and relation triple extraction (Jin et al., 2021; Nie et al., 2024; Li et al., 2024b). However, as discussed earlier, domain-specific unstructured data sources are inherently limited in scope and coverage, constraining the effectiveness of these text-based extraction methods. Moreover, the specialized terminology and complex linguistic patterns in domain-specific texts further exacerbate the challenges of accurate knowledge extraction. ② The latter category focuses on mining implicit knowledge from the existing graph itself, typically via knowledge graph completion (Wei et al., 2024), which predicts missing triples through internal inference within the existing KG. While these solutions can discover implicit relationships, the knowledge they generate is fundamentally bounded by the original KG’s scope and cannot introduce genuinely new external knowledge. Given the limitations of this existing research and the rich, continuously updated knowledge available in comprehensive GKGs like Wikipedia and YAGO, a natural question arises: Can we systematically leverage the abundant knowledge in GKGs to supplement and enrich DKGs?
@@
To this end, we propose a new and important task setting, i.e., domain-specific knowledge graph fusion (DKGF), which aims to mine the highly relevant knowledge in the general knowledge graph to enhance the knowledge coverage of domain-specific knowledge graph, as illustrated in Figure 1. Unlike the widely studied knowledge graph alignment task (Zeng et al., 2020; Ge et al., 2021; Wu et al., 2023; Su et al., 2025; Tian et al., 2024; Ding et al., 2025), which focuses on identifying entities that are simultaneously present and semantically equivalent across different KGs, the DKGF task essentially focuses on how to extract knowledge from a large-scale semantically complex GKG that is closely related to the target domain but have not yet been included in the DKG, and then effectively integrate them, while maintaining the semantic consistency and structural integrity of the DKG itself. As a result, the DKGF task poses new challenges:
@@
Challenge I: High Ambiguity of Domain Relevance. Firstly, GKGs contain a large amount of knowledge with low similarity to that in DKGs, yet such information may still provide critical factual associations for enriching DKGs in reality. As illustrated in Figure 1, during the mining process of relevant knowledge from GKGs, the academic institution “Brookings Institution” from Washington D.C. lacks explicit connections to entities in the political crisis event DKG (e.g., “Iran”), resulting in low semantic similarity between entities. Nonetheless, in reality, “Brookings Institution” maintains many important connections with DKG entities and is highly likely to constitute factual associations with them. This discrepancy creates ambiguity in determining whether knowledge from GKGs is truly domain-relevant, making it difficult to directly apply traditional graph matching methods or language models (Pan et al., 2024a) to this new task.
@@
To tackle the root causes of DKGF, we propose ExeFuse, a neuro-symbolic framework that reframes fusion not as static matching, but as a Fact-as-Program process where knowledge facts are modeled as executable programs transformable against target domain logic. Specifically, to resolve the high ambiguity of domain relevance, ExeFuse introduces neuro-symbolic program execution, which maps discrete GKG facts into a continuous space and treats logic rules as transition operators to infer logical reachability beyond surface dissimilarity. Simultaneously, to address cross-domain knowledge granularity misalignment, we employ a target space grounding mechanism to calibrate knowledge specificity. This approach constructs a target state space and utilizes an executability core to verify that the transformed knowledge successfully grounds into the valid semantic region of the DKG, ensuring that the fused facts are not only logically relevant but also compatible with the specific granularity and structural patterns of the target domain. Notably, to fill in the gap of DKGF benchmarks, we construct two new datasets: DKGF(W-I) and DKGF(Y-I), and offer 21 benchmark configurations for systematic evaluation. Extensive experiments highlight the value of the new task and demonstrate the effectiveness of ExeFuse, providing the first standardized evaluation suite for this emerging task.
@@
    We define a new task setting, domain-specific knowledge graph fusion (DKGF), highlighting its distinct requirements compared to traditional tasks.
@@
    We propose the Fact-as-Program paradigm, a neuro-symbolic approach that models fact fusion as a verifiable logical execution process in a quasi-symbolic space.
@@
    We provide theoretical analysis to guarantee the expressiveness of our neuro-symbolic mapping and the validity of the granularity calibration.
  • $\bullet$
    We develop two new benchmark datasets to establish a standardized evaluation for the DKGF task. Comprehensive experiments demonstrate the effectiveness, efficiency, and generalization of our ExeFuse framework.
@@ 2. Related work
To the best of our knowledge, DKGF is a novel task in domain-specific data integration. Closely related research can be categorized into extraction from external unstructured sources, internal knowledge reasoning, and cross-graph entity alignment.
Extraction from External Unstructured Sources.  This category focuses on acquiring structured knowledge from external unstructured texts to enrich KGs.
  • Relation Triple Extraction (RTE) aims to extract structured triplets from unstructured text (Hu et al., 2023; Li et al., 2024a; Yang et al., 2025b). It has evolved from supervised pipelines (Wang and Lu, 2020; Ye et al., 2022a; Lin et al., 2019) to joint extraction models. To address data scarcity in specialized domains, research has shifted towards zero-shot settings (Li et al., 2025c). Key approaches include data augmentation methods (Chia et al., 2022; Guo et al., 2024; Gong and Eldardiry, 2024) and knowledge-driven methods (Kim et al., 2023). Recently, bilevel optimization (Li et al., 2025b) and LLM-based extraction (Yang et al., 2025b) have been introduced to enhance generalization.
  • Entity Linking (EL) identifies entity mentions in text and links them to corresponding entities in KGs. While early works relied on handcrafted features (Graciotti et al., 2025), the field has moved towards dense retrieval and generative paradigms (van Hulst et al., 2020). Bi-encoder architectures (e.g., BLINK (Wu et al., 2020)) and generative EL methods (e.g., GENRE (Cao et al., 2021)) are widely used, with recent LLM-augmented pipelines (Graciotti et al., 2025; Ding et al., 2024) leveraging in-context learning.
However, both RTE and EL heavily rely on the availability and quality of domain-specific corpora, which are often scarce or noisy in specialized domains.
Internal Knowledge Reasoning.  Unlike extraction methods that seek external information, this category focuses on inferring missing facts solely based on existing information within the KG. Research in this area primarily centers on knowledge graph completion (KGC), which predicts missing triples within a single KG. Embedding-based techniques (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016; Sun et al., 2019; Yao et al., 2019a) project elements into continuous vector spaces to score plausibility. PLM-based methods (Yao et al., 2019a; Kim et al., 2020; Wang et al., 2021; Lv et al., 2022a; Devlin et al., 2019b) and generative approaches (Saxena et al., 2022; Chen et al., 2022a; Ye et al., 2022b) further enhance performance. Although effective for link prediction, KGC is inherently limited to the closed-world assumption of the existing graph and cannot discover new entities or facts from the outside world.
Entity Alignment.  Entity alignment (EA) aims to identify equivalent entities across KGs. Early translation-based methods (Zhao et al., 2025b; Sun et al., 2018; Bordes et al., 2013; Chen et al., 2017) utilize distance models but often struggle with complex structures. GNN-based approaches (Wang et al., 2018; Chen et al., 2022b; Mao et al., 2021; Xu et al., 2021; Cai et al., 2022; Liu et al., 2023; Zeng et al., 2024) leverage graph convolutions to capture neighborhood topologies, while recent solutions integrate side information (Wang et al., 2023a; Jiang et al., 2024b; Tang et al., 2020) or leverage LLMs for few-shot semantic matching (Chen et al., 2024b; Jiang et al., 2024a; Zhao et al., 2025a; Yang et al., 2024).
@@
Task Distinction and Gap Analysis.  Despite these efforts, existing research predominantly relies on extracting knowledge from unstructured data or completing KGs through internal reasoning, while the scope and quality of knowledge integration remain significantly limited (Yu et al., 2022; Li et al., 2025a). This situation exposes a critical gap: There has been little systematic exploration of how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. DKGF is therefore an essential and indispensable task. Furthermore, this new task differs fundamentally from existing ones:
  • Vs. RTE: RTE operates on unstructured text, whereas DKGF leverages structured GKG.
  • Vs. EL: EL resolves text-to-KG disambiguation, whereas DKGF enhances DKG coverage via cross structured graph discovery and fusion.
  • Vs. KGC: KGC focuses on reasoning within a single KG, while DKGF integrates knowledge from external GKG.
  • Vs. EA: EA outputs same entity pairs across KGs, whereas DKGF identifies missing but relevant entities and produces new facts.
Consequently, existing research cannot address the core challenges of DKGF: high ambiguity of domain relevance and cross-domain granularity misalignment. While neuro-symbolic reasoning (Alshahrani et al., 2017; Cheng et al., 2024; Spillo et al., 2024; Liu et al., 2025; Cheng and Sun, 2024) provides the explicit logical grounding needed to resolve such ambiguity, it has typically been confined to single-graph inference. Inspired by this, our ExeFuse adopts a Fact-as-Program paradigm, extending neuro-symbolic execution to dynamically validate logical relevance and granularity beyond surface matching.